OpenStudy (anonymous):

Suppose y varies directly with x and y=-1/2 when x=2/3. Write an equation relating x and y. A. y=4/3x B. y=-4/3x C. y=-3/4x D. y+3/4x

OpenStudy (radar):

-1/2=k2/3 k=-3/4 y=(-3/4)x C.